OpenStudy (anonymous):

a cone shape container has a base of 113.04cubic centimeter and a height of 20cm. if it is full, how much water can it hold?

OpenStudy (anonymous):

\[V = \frac{ 1 }{ 3 }\Pi r ^{2}h\]

OpenStudy (anonymous):

pi r^2 = base = 113.04

OpenStudy (anonymous):

V = 1/3 * 113.04 * 20

OpenStudy (anonymous):

how did the base become cubic centimeter?

OpenStudy (anonymous):

base should be centimeter square

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@lambchamps is my approach is right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

why did you substitute the 113.04 in pi r^2?

OpenStudy (anonymous):

base of the cone is circle and area of circle is pi r^2

OpenStudy (anonymous):

have you understand @LOL10

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

is the answer 753.6??

OpenStudy (anonymous):

what will be the unit?

OpenStudy (anonymous):

yes you are right qed

OpenStudy (anonymous):

@uzumakhi what will be the unit?

OpenStudy (anonymous):

\[cm ^{3}\]

OpenStudy (anonymous):

thank you